# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ A.2 Magnus Expansion of the Transmon Hamiltonian
To evaluate the integrals in Eq. A.5-A.7, it is most convenient to decompose $\mathcal{H}_{\text{mov}}$ in terms of Fourier components
$\mathcal{H}_{\text{mov}}(t)=\sum_{m}h_{m}e^{im\omega t}\,,$ (A.10)
@@
$h_{0}=\mathcal{H}^{0}_{\widehat{n}}-E_{J}\sum_{k}\frac{F_{0,k}e^{-i{\widehat{\varphi}}_{k}}+F_{-0,k}^{*}e^{i{\widehat{\varphi}}_{k}}}{2}\,,$ (A.11)
$h_{m\neq 0}=-E_{J}\sum_{k}\frac{F_{m,k}e^{-i{\widehat{\varphi}}_{k}}+F_{-m,k}^{*}e^{i{\widehat{\varphi}}_{k}}}{2}\,,$ (A.12)
@@
$e^{i\theta_{k}(t)}=\sum_{m}e^{im\omega t}F_{m,k}\,.$ (A.13)
The explicit expression for $F_{m,k}$ is given in Eq. 24 in the main text.
@@
$\mathcal{H}_{\text{eff}}^{1}=\sum_{m,n}I^{(1)}_{mn}[h_{m},h_{n}]=\sum_{m\neq 0}\left(I^{(1)}_{m0}-I^{(1)}_{0m}\right)[h_{m},h_{0}]\,,$ (A.15)
@@
$[e^{i{\widehat{\varphi}}_{k}},G(\hat{n}_{k})]$ $=-e^{i{\widehat{\varphi}}_{k}}\Delta_{k}G(\hat{n}_{k})\,,$ (A.17)
$[e^{-i{\widehat{\varphi}}_{k}},G(\hat{n}_{k})]$ $=\Delta_{k}G(\hat{n}_{k})e^{-i{\widehat{\varphi}}_{k}}\,,$ (A.18)
@@
$\mathcal{H}^{2}_{\text{eff}}=\sum_{mn}[h_{m},[h_{n},h_{0}]]\left(I^{(2)}_{mn0}+I^{(2)}_{0nm}-I^{(2)}_{m0n}-I^{(2)}_{n0m}\right)\,.$ (A.23)
@@
$H^{2}_{B}$ $=$ $\sum_{n\neq 0}[h_{0}-\mathcal{H}^{0}_{\widehat{n}},[h_{n},h_{0}]]\left(2I^{(2)}_{0n0}-I^{(2)}_{00n}-I^{(2)}_{n00}\right)\,,$ (A.24b)
$H^{2}_{C}$ $=$ $\sum_{n\neq 0}[\mathcal{H}^{0}_{\widehat{n}},[h_{n},h_{0}]]\left(2I^{(2)}_{0n0}-I^{(2)}_{00n}-I^{(2)}_{n00}\right)\,.$ (A.24c)
@@
$\left(I^{(2)}_{mn0}+I^{(2)}_{0nm}-I^{(2)}_{m0n}-I^{(2)}_{n0m}\right)_{S}$ $=$ $\frac{1}{2mn\omega^{2}},\quad(m,n\neq 0)\,,$
@@
In Eq. A.25, we have symmetrized the result with respect to $m\leftrightarrow n$ (indicated by the subscript $S$), because $m,n$ can be exchanged in Eq. A.24 due to $[h_{m},h_{n}]=0$. The final step is to compute the commutators, and we obtain Eqs. (29) in the main text.
Appendix B Role of Hilbert-space Dimension and Initial State on Freezing
Figure 14: Stroboscopic time evolution of normalized $\widehat{N}(t)$ for different driving strengths for the square drive in Eq. 3a. The system size is $L=6$, with an on-site Hilbert space dimension $\mathcal{H}_{D}=4$. All other parameters are as in Fig. 2.
In this appendix, we comment on the effect of increasing local Hilbert-space dimension (${\cal{H}}_{D}$) and the choice of initial state (e.g. picked from the eigenspectrum) on the robustness of freezing. Consider for simplicity starting with ${\cal{H}}_{D}=2$ (i.e. equivalent to a spin1/212-1/2- 1 / 2 system). Recall that for the square-wave drive within the zeroth-order Magnus expansion, the non-integrable portion of the Hamiltonian is:
$\dfrac{\omega E_{J}}{A\pi}\sin\Big{(}\dfrac{\pi A}{\omega}\Big{)}\sum_{k}\cos{\hat{\varphi}_{k}}-\dfrac{\omega E_{J}}{A\pi}\Big{(}1-\cos\Big{(}\dfrac{\pi A}{\omega}\Big{)}\Big{)}\sum_{k}\sin{\hat{\varphi}_{k}}.$ (B.26)
The matrix elements in the basis states of interest are
$\langle m_{k}|\cos{\hat{\varphi}_{k}}|n_{k}\rangle,\ \langle m_{k}|\sin{\hat{\varphi}_{k}}|n_{k}\rangle,$ (B.27)
@@ Appendix B Role of Hilbert-space Dimension and Initial State on Freezing
If we recall the condition for freezing under the square-wave drive, we find that $A/\omega=2\mathbb{Z}$, where both terms in Eq. B.26 vanish. However, in Fig. 12(a), we notice additional peaks not only at $A/\omega=2\mathbb{Z}$ (even integer multiples) but also at odd integer values. For odd integer ratios of $A/\omega$, the first term in Eq. B.26 goes to zero. Although the second term remains non-zero, its impact is less pronounced because it involves off-diagonal elements. In Fig. 12(b), we observe that as the ratio of amplitude to frequency increases, ${\cal{F}}$ approaches $1$ regardless of whether we are near or away from freezing. This is due to the effect of the nonintegrable part of the effective Hamiltonian being diminished quantitatively. It is noteworthy that a similar trend is evident in the plot of the Lyapunov exponent shown in Fig. 7(a). In Fig. 12(c) we show the initial state dependence on ${\cal{F}}$ as a function of $A/\omega$, which demonstrates that the robustness associated with the freezing peaks for the higher excited states (see Fig. 13) diminishes (Eq. B.26). We only observe the remaining freezing peaks for ${\cal{F}}$ at the even integer ratios of $A/\omega$. Finally, we have also studied the fate of freezing for the same square-wave drive protocol with local Hilbert-space dimension, ${\cal{H}}_{D}=4$, as shown in Fig. 14.

### v2

@@ B.2 Magnus Expansion of the Transmon Hamiltonian
To evaluate the integrals in Eq. (S.7)-(S.9), it is most convenient to decompose $\mathcal{H}_{\text{mov}}$ in terms of Fourier components
$\mathcal{H}_{\text{mov}}(t)=\sum_{m}h_{m}e^{im\omega t}\,,$ (S.12)
@@
$h_{0}=\mathcal{H}^{0}_{\widehat{n}}-E_{J}\sum_{k}\frac{F_{0,k}e^{-i{\widehat{\varphi}}_{k}}+F_{-0,k}^{*}e^{i{\widehat{\varphi}}_{k}}}{2}\,,$ (S.13)
$h_{m\neq 0}=-E_{J}\sum_{k}\frac{F_{m,k}e^{-i{\widehat{\varphi}}_{k}}+F_{-m,k}^{*}e^{i{\widehat{\varphi}}_{k}}}{2}\,,$ (S.14)
@@
$e^{i\theta_{k}(t)}=\sum_{m}e^{im\omega t}F_{m,k}\,.$ (S.15)
The explicit expression for $F_{m,k}$ is given in Eq. (13) in the main text.
@@
$\mathcal{H}_{\text{eff}}^{1}=\sum_{m,n}I^{(1)}_{mn}[h_{m},h_{n}]=\sum_{m\neq 0}\left(I^{(1)}_{m0}-I^{(1)}_{0m}\right)[h_{m},h_{0}]\,,$ (S.17)
@@
$[e^{i{\widehat{\varphi}}_{k}},G(\hat{n}_{k})]$ $=-e^{i{\widehat{\varphi}}_{k}}\Delta_{k}G(\hat{n}_{k})\,,$ (S.19)
$[e^{-i{\widehat{\varphi}}_{k}},G(\hat{n}_{k})]$ $=\Delta_{k}G(\hat{n}_{k})e^{-i{\widehat{\varphi}}_{k}}\,,$ (S.20)
@@
$\mathcal{H}^{2}_{\text{eff}}=\sum_{mn}[h_{m},[h_{n},h_{0}]]\left(I^{(2)}_{mn0}+I^{(2)}_{0nm}-I^{(2)}_{m0n}-I^{(2)}_{n0m}\right)\,.$ (S.25)
@@
$H^{2}_{B}$ $=$ $\sum_{n\neq 0}[h_{0}-\mathcal{H}^{0}_{\widehat{n}},[h_{n},h_{0}]]\left(2I^{(2)}_{0n0}-I^{(2)}_{00n}-I^{(2)}_{n00}\right)\,,$ (S.26b)
$H^{2}_{C}$ $=$ $\sum_{n\neq 0}[\mathcal{H}^{0}_{\widehat{n}},[h_{n},h_{0}]]\left(2I^{(2)}_{0n0}-I^{(2)}_{00n}-I^{(2)}_{n00}\right)\,.$ (S.26c)
@@
$\left(I^{(2)}_{mn0}+I^{(2)}_{0nm}-I^{(2)}_{m0n}-I^{(2)}_{n0m}\right)_{S}$ $=$ $\frac{1}{2mn\omega^{2}}-\frac{1}{2m^{2}\omega^{2}}\delta_{m,-n},\quad(m,n\neq 0)\,,$
@@
In Eq. (S.27), we have symmetrized the result with respect to $m\leftrightarrow n$ (indicated by the subscript $S$), because $m,n$ can be exchanged in Eq. (S.26) due to $[h_{m},h_{n}]=0$. The final step is to compute the commutators, and we obtain Eqs. (18) in the main text.
Appendix C Discussion of Secondary Freezing Peaks
Here we comment on some additional details associated with the freezing phenomenology. We seek to explain both the emergent conservation of $\hat{N}$ at discrete points away from freezing and how the choice of initial state affects the robustness of freezing.
Consider for simplicity starting with ${\cal{H}}_{D}=2$ (i.e. equivalent to a spin1/212-1/2- 1 / 2 system). Recall that for the square-wave drive within the zeroth-order Magnus expansion, the non-integrable portion of the Hamiltonian is:
$\dfrac{\omega E_{J}}{A\pi}\sin\Big{(}\dfrac{\pi A}{\omega}\Big{)}\sum_{k}\cos{\hat{\varphi}_{k}}-\dfrac{\omega E_{J}}{A\pi}\Big{(}1-\cos\Big{(}\dfrac{\pi A}{\omega}\Big{)}\Big{)}\sum_{k}\sin{\hat{\varphi}_{k}}.$ (S.28)
The matrix elements in the basis states (Fock states) of interest are
$\langle m_{k}|\cos{\hat{\varphi}_{k}}|n_{k}\rangle,\ \langle m_{k}|\sin{\hat{\varphi}_{k}}|n_{k}\rangle,$ (S.29)
@@ Appendix C Discussion of Secondary Freezing Peaks
If we recall the condition for freezing under the square-wave drive, we find that $A/\omega=2\mathbb{Z}$, where both terms in Eq. (S.28) vanish. However, in Fig. (LABEL:fig:Hilbertdepen)(a), we notice additional peaks not only at $A/\omega=2\mathbb{Z}$ (even integer multiples) but also at odd integer values. For odd integer ratios of $A/\omega$, the first term in Eq. (S.28) goes to zero. Although the second term remains non-zero, its impact is less pronounced because it involves off-diagonal elements. In Fig. LABEL:fig:Hilbertdepen(b), we observe that as the ratio of amplitude to frequency increases, ${\cal{F}}$ approaches $1$ regardless of whether we are near or away from freezing. This is due to the effect of the nonintegrable part of the effective Hamiltonian being diminished quantitatively. It is noteworthy that a similar trend is evident in the plot of the Lyapunov exponent shown in Fig. LABEL:fig:classicalLE(a). In Fig. LABEL:fig:Hilbertdepen(c) we show the initial state dependence on ${\cal{F}}$ as a function of $A/\omega$, which demonstrates that the robustness associated with the freezing peaks for the higher excited states (see Fig. LABEL:fig:nexptH(0)) diminishes (Eq. (S.28)). We only observe the remaining freezing peaks for ${\cal{F}}$ at the even integer ratios of $A/\omega$.
Appendix D Dependence of Semi-Classical Freezing on $\omega$
Here we comment on the threshold on $\omega$ below which semi-classical freezing vanishes. In Fig. LABEL:fig:omega_threshold_vs_alpha(a) we plot the MLE for a one-dimensional chain of length $8$ due to the square-wave drive, as a function of the ratio between drive frequency $\omega$ and the transmon frequency $\nu$. The curves show the first six freezing points (as determined by the zeroth-order Magnus expansion) given by Eq. (40a). We find that the $\omega$-threshold has a clear dependence on the value of $\alpha=A/\omega$ indicating that the $\omega$-threshold is itself a function of the drive amplitude $A$. Interestingly, we observe that the $\omega$-threshold seems to converge towards an $O(\nu)$ number for increasingly large values of $\alpha$.
In Fig. LABEL:fig:omega_threshold_vs_alpha(b) we plot the MLE for the same chain and drive, as a function of the drive amplitude $A$ at fixed frequencies $\omega/(2\pi)=15$ GHz and $\omega/(2\pi)=9.32$ GHz. We observe that the first freezing point at $\alpha=2$ that is present in the former case, does not appear for the smaller drive frequency. The remaining freezing points are present for both frequencies. This is in alignment with Fig. LABEL:fig:omega_threshold_vs_alpha(a) which indicates that the $\omega$ threshold for the first freezing point lies at around $\omega/(2\pi)\sim 11$ GHz.
Supplemental Figure 1. Freezing for large local Hilbert space dimension:
(a) Time averaged expectation of $N(t)/N(0)$ is plotted as a function of $A/\omega$ for a linear chain of $4$ transmons and local Hilbert space dimension $8$ for the square wave drive. All of the other parameters are as in Fig. LABEL:fig:ndtstrobo in the main text. Note that there is a slight shift of the freezing point due to higher order corrections in the Magnus expansion. (b) Evolution of $\hat{N}(t)$ for the same system at and away from freezing. Note that we use the renormalized freezing point, $A/\omega=2.1$, determined numerically from (a).
Supplemental Figure 2: Dependency of freezing on local Hilbert space dimension and system size:
(a) The time averaged (over 100 Floquet cycles) expectation of $N(t)/N(0)$ as function of $A/\omega$ for different system size ($L$) keeping the Hilbert space dimension ($\cal{D}$) fixed. (b) Same quantity for different $\cal{D}$ for fixed $L$. In both cases the $\cos(\phi)$ nonlinearity are kept up to order 8. As an initial state we chose the ground state of $H(0)$. All of the other parameters are as in Fig. LABEL:fig:ndtstrobo in the main text.
Supplemental Figure 3: Behavior of secondary freezing points:
(a) ${\cal{F}}$ (Eq. (6)) averaged over 100 Floquet cycles as a function of $\alpha=A/\omega$ for two different system sizes ($L=6,8$) and onsite Hilbert space (${\cal{H}}_{D}=2$), with |$\ket{\psi(0)}$ chosen as ground state of $H(0)$. In addition to the sharp peaks at even integer $\alpha$, peaks at odd integers are also seen. (b) Same as in (a), but with ${\cal{H}}_{D}=3$. With increasing $\alpha$, $\cal{F}$ flattens out since $\mathcal{H}^{0}_{\widehat{\varphi}}$ weakens compared to $\mathcal{H}^{0}_{\widehat{n}}+f(t)\mathcal{H}^{\rm{drive}}_{\widehat{n}}$. Furthermore, as anticipate, we see a reduction in the extent of the conservation of $\hat{N}$ away from freezing for the larger system size. (c) For $L=8$ transmons and ${\cal{H}}_{D}=3$, variation of ${\cal{F}}$ for different |$\ket{\psi(0)}$, drawn from the eigenspectrum of $H(0)$ in Fig. LABEL:fig:nexptH(0). All other parameters are as in Fig. LABEL:fig:ndtstrobo in the main text.
Supplemental Figure 4: $\hat{\textbf{N}}$ expectation of initial states
The expectation value of the operator $\hat{N}$ is calculated with respect to the eigenstates of the Hamiltonian $H(0)$ and plotted against the corresponding eigen-energy for square wave drive. The eigenstates marked by arrows are used as |$\ket{\psi(0)}$ to calculate the observables in Eq. (6) and Fig. LABEL:fig:Hilbertdepen.
Supplemental Figure 5: Frequency threshold of semi-classical freezing:
(a) MLE plotted as a function of $\omega/\nu$ at various zeroth-order freezing points of the square-wave drive. The system is a length $8$ chain initialized as $1010\ldots$. For larger freezing points, $\alpha=A/\omega$, the omega threshold, below which freezing does not occur, decreases while appearing to approach an $O(\nu)$ number. (b) MLE plotted as a function of $\alpha$ at fixed frequencies $\omega/(2\pi)=15$ GHz and $\omega/(2\pi)=9.32$ GHz of the square-wave drive. Notably, the $\alpha=2$ freezing point is absent in the latter case indicating that the omega threshold for the freezing point lies between $9.32$ and $15$ GHz.